OpenStudy (anonymous):

How to type a 3x3 determinant??

OpenStudy (unklerhaukus):

``` \[ \det{\bf A}= \left| \begin{array}{ccc} a_{11}&a_{12}&a_{13}\\ a_{21}&a_{22}&a_{23}\\ a_{31}&a_{32}&a_{33} \end{array} \right| \] ``` \[ \det{\bf A}= \left| \begin{array}{ccc} a_{11}&a_{12}&a_{13}\\ a_{21}&a_{22}&a_{23}\\ a_{31}&a_{32}&a_{33} \end{array} \right| \]

OpenStudy (anonymous):

Nice!! Thanks!! How do you type that grey box?

OpenStudy (unklerhaukus):

``` like this ```

OpenStudy (anonymous):

It doesn't work? :\

OpenStudy (unklerhaukus):

no space before the the `'s on the third line

OpenStudy (anonymous):

`Testing ` Oh! It works!! Thanks again!!!

OpenStudy (unklerhaukus):

``` \checkmark ```

OpenStudy (anonymous):

` How come your box is soooooo long in terms of its length!!! `

OpenStudy (anonymous):

\[\checkmark\] o.O!